# VIST3A: Text-to-3D by Stitching a Multi-view Reconstruction Network to a Video Generator

### v1

@@ Abstract
The rapid progress of large, pretrained models for both visual content generation and 3D reconstruction opens up new possibilities for text-to-3D generation. Intuitively, one could obtain a formidable 3D scene generator if one were able to combine the power of a modern latent text-to-video model as “generator” with the geometric abilities of a recent (feedforward) 3D reconstruction system as “decoder”. We introduce VIST3A, a general framework that does just that, addressing two main challenges. First, the two components must be joined in a way that preserves the rich knowledge encoded in their weights. We revisit model stitching, i.e., we identify the layer in the 3D decoder that best matches the latent representation produced by the text-to-video generator and stitch the two parts together. That operation requires only a small dataset and no labels. Second, the text-to-video generator must be aligned with the stitched 3D decoder, to ensure that the generated latents are decodable into consistent, perceptually convincing 3D scene geometry. To that end, we adapt direct reward finetuning, a popular technique for human preference alignment. We evaluate the proposed VIST3A approach with different video generators and 3D reconstruction models. All tested pairings markedly improve over prior text-to-3D models that output Gaussian splats. Moreover, by choosing a suitable 3D base model, VIST3A also enables high-quality text-to-pointmap generation.
@@
footnotetext: Project page: https://gohyojun15.github.io/VIST3A/
@@ 1 Introduction
With image and video generators now a commodity, text-to-3D models that produce 3D scenes from text prompts have become a new research frontier, with applications in AR/VR, gaming, robotics, and simulation. Early methods for 3D generation adopt Score Distillation Sampling (SDS) (Poole et al., 2023; Tang et al., 2024b; Wang et al., 2023b; Chen et al., 2024b) to optimize a 3D representation, e.g. a NeRF (Mildenhall et al., 2021; Müller et al., 2022) or 3D Gaussian Splats (3DGS, Kerbl et al., 2023) under a pretrained 2D diffusion prior (Rombach et al., 2022). A drawback these methods have in common is the need for slow per-scene optimization. Another line of work uses multi-stage pipelines that first synthesize images and then lift them to 3D with a separate model (Tang et al., 2024a; Xu et al., 2024b; Zhang et al., 2024b) or with per-scene optimization (Gao et al., 2024; Wu et al., 2024a; Yu et al., 2024b); employ progressive warping and refinement (Shriram et al., 2025; Yu et al., 2025; 2024a); or sequentially chain multiple generative modules (Yang et al., 2025b; Engstler et al., 2025). The multi-stage design not only increases model complexity and engineering effort, but also makes such models prone to error accumulation (Lin et al., 2025; Meng et al., 2025).
@@
A recent trend is to directly generate the 3D representation with end-to-end latent diffusion models  (LDMs, Schwarz et al., 2025; Lan et al., 2024; Li et al., 2025b; a; Bahmani et al., 2025). A prominent line of work starts from pretrained 2D image (Esser et al., 2024; Rombach et al., 2022) or video models (Genmo Team, 2024; Yang et al., 2024b) and finetunes them to output multi-view 2D latents, reusing the pretrained priors (Szymanowicz et al., 2025; Liang et al., 2025; Schwarz et al., 2025; Lin et al., 2025; Yang et al., 2025c; Go et al., 2025a; b). Subsequently, a VAE-style decoder is trained to decode those latents into the desired 3D representation, see Fig. 2. The LDM-like design unifies 2D generation and multi-view reconstruction within the latent space and enables efficient 3D scene generation with a compact, well-amortized decoder.
@@
Still, two key limitations remain. First, we argue that the Achilles heel of existing 2D-to-3D diffusion models is the decoder. By simply repurposing the 2D VAE to produce 3D outputs, the network must learn 3D reconstruction more or less from scratch, which requires extensive training and large datasets that are hard to obtain (Yang et al., 2025c; Szymanowicz et al., 2025; Go et al., 2025b). This practice becomes increasingly problematic as new, better 3D foundation models emerge  (Wang et al., 2025d; a; 2024b; Zhang et al., 2025) and the ad-hoc trained decoders of text-to-3D models fall further behind the state of the art in 3D vision.
@@
Second, the prevalent training scheme tends to suffer from weak alignment between the generative model and the VAE decoder. Typically, the former is finetuned on multi-view datasets with a generative objective like a diffusion loss (Song et al., 2020; Sohl-Dickstein et al., 2015; Ho et al., 2020) or flow matching (Liu et al., 2023; Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023), which only indirectly promotes 3D-consistent latents. Moreover, the separate training may cause the latents, even if 3D-consistent, to be out of domain from the perspective of the decoder. To mitigate that misalignment, it has been proposed to add rendering losses that promote decodable latents (Lin et al., 2025). However, the resulting objective is based on single-step sampling and does not sufficiently take into account the denoising trajectory, leading to weak alignment at inference.
@@
We introduce VIST3A: VIdeo VAE STitching and 3D Alignment. The proposed method consists of two complementary components that address the above-mentioned limitations, see Fig. 2. First, we resort to the concept of model stitching (Pan et al., 2023; Lenc & Vedaldi, 2015; Bansal et al., 2021; Csiszárik et al., 2021; Yang et al., 2022) to leverage powerful, pretrained feedforward 3D models for decoding, rather than start from scratch. The idea is to attach the relevant part of a 3D reconstruction network as a “decoder” to the latent space of a video VAE. For this to work, there need to be one or more layers in the 3D model whose activations are similar (up to a linear transformation) to those in the VAE’s latent space, despite their independent pretraining. Perhaps surprisingly, this turns out to be the case. For the 3D model, we identify the layer with the most linear relation to the LDM latents, slice the network before that layer, and retain the downstream portion as 3D decoder. After fitting a single, linear stitching layer (in closed form), the VAE latent space already matches the expected input of the 3D decoder well, such that subsequent fine-tuning will be minor and not degrade the respective generative and 3D reasoning capabilities of the two base models.
@@
Second, we further improve alignment between the generative model and the stitched decoder through direct reward finetuning (Clark et al., 2023; Xu et al., 2023; Prabhudesai et al., 2024; Wu et al., 2024c; Shen et al., 2025). In that technique, commonly used to align diffusion models with human preferences, reward signals are defined based on the “goodness” of the VAE output – in our setting, the visual quality and 3D consistency of the decoded 3D representations. Maximizing these rewards encourages the LDM to produce latents that are 3D-consistent and lie within the decoder’s input domain, ensuring high-quality outputs. Importantly, our alignment compares video model outputs and images rendered from the generated 3D scenes, hence it does not require labels.
@@ 2 Related Works
3D generation. Recent works have explored various 3D representations for generative modelling, including point clouds (Mo et al., 2023; Nichol et al., 2022; Vahdat et al., 2022), meshes (Xu et al., 2024a; Woo et al., 2024), voxel grids (Sanghi et al., 2023), NeRFs (Chen et al., 2023; Müller et al., 2022; Mildenhall et al., 2021), and 3DGS (Henderson et al., 2024; Zhang et al., 2024a; Kerbl et al., 2023). Score distillation using 2D diffusion models is time-consuming, as it requires per-scene test time optimization (Wang et al., 2023a; Shi et al., 2023; Wang et al., 2023b), while multi-stage pipelines (Yu et al., 2024b; Liu et al., 2024; Zheng et al., 2025) lack robustness and create significant engineering overhead. For further details on multi-stage pipelines, please refer to Appendix A.
@@
To tackle such situations, recent methods (Szymanowicz et al., 2025; Liang et al., 2025; Schwarz et al., 2025; Lin et al., 2025; Yang et al., 2025c; Go et al., 2025a; b; Bahmani et al., 2025) repurpose the comprehensive knowledge of the visual world that is implicit in 2D image generators. The general strategy is to finetune a pretrained 2D model on multi-view data, by using generative losses to enforce cross-view consistency. In many cases training is further supported by additional 3D cues like camera poses (Li et al., 2024; Go et al., 2025b), depthmaps (Go et al., 2025a; Yang et al., 2025c), or pointmaps (Szymanowicz et al., 2025). The resulting multi-view latents are decoded to 3D scenes with a dedicated VAE-style decoder, meaning that 3D reasoning capabilities must be rebuilt from scratch, and that they are only weakly aligned with the generator output – limitations which we address with VIST3A.
@@
Learned 3D reconstruction. A notable trend in 3D computer vision is the trend to move away from multi-stage pipelines and iterative optimization towards end-to-end, feedforward 3D modelling. Classical reconstruction pipelines based on SfM (Hartley & Zisserman, 2003; Schönberger & Frahm, 2016) and MVS (Furukawa et al., 2015; Schönberger et al., 2016) require incremental, iterative optimization, whereas recent advances like DUSt3R (Wang et al., 2024b) and MASt3R (Leroy et al., 2024) directly predict 3D point maps in one forward pass. Several follow-up works have further reduced test-time optimization (Tang et al., 2025; Wang et al., 2025b; Yang et al., 2025a). Likewise, 3D Gaussian splatting has evolved from per-scene optimization to feedforward prediction (Charatan et al., 2024; Chen et al., 2024a; Ye et al., 2024). Once more, data scaling has been a critical factor (Wang et al., 2025a; d). Consequently, replicating the 3D capabilities of recent feedforward models as part of VAE training would be difficult and costly. VIST3A offers a solution by reusing, rather than rebuilding, models like AnySplat (Jiang et al., 2025), VGGT (Wang et al., 2025a), or MVDUSt3R (Tang et al., 2025).
@@
Model stitching. Recomposing the heads and tails of two different networks was initially studied as a way to assess the equivariance of neural representations (Lenc & Vedaldi, 2015), and as an experimental tool to compare two different representations (Csiszárik et al., 2021; Bansal et al., 2021). To ensure invariance against trivial affine transformations, the head of some trained network $A$ is normally attached to the tail of another network $B$ via a linear, trainable stitching layer. Besides revealing similarities between networks that common metrics like CKA (Kornblith et al., 2019) would miss, it was also found that different architectures that were trained on the same data can often be stitched into a new, hybrid model with minimal degradation (Bansal et al., 2021). This has opened the door for practical uses of stitching, e.g. DeRy (Yang et al., 2022) for resource-constrained reassembly of pretrained models and SN-Net (Pan et al., 2023) to build networks with varying scales. Going one step further, we demonstrate that strong 3D VAEs [^1] To be consistent with existing literature (Lan et al., 2024; Yang et al., 2025c), we also use the term “3D VAE”, although the mapping from 2D images to 3D scene is, technically, not a variational auto-encoder. can be obtained by stitching a foundational 3D model to the latent space of a video VAE as its decoder, even if they were trained independently on different data.
@@ 3.1 Model Stitching for 3D VAE Construction
Step 1: Finding the stitching index and initialization. To identify the layer $k^{*}$ in the 3D model whose representation is most compatible with the VAE latent, we first push a set of $N$ samples through the encoder $\mathcal{E}$ to obtain their latents $\mathbf{B}\in\mathbb{R}^{N\times D_{\mathcal{E}}}$. Then, we scan over candidate layers $k\in\{1,...,l-1\}$ of the 3D model and, for each layer in turn, extract the activations $\mathbf{A}_{k}\in\mathbb{R}^{N\times D_{F}}$ and fit the linear stitching layer $\mathbf{S^{*}}_{k}\in\mathbb{R}^{D_{\mathcal{E}}\times D_{F}}$ that best recovers the activations of the 3D model at layer $k$, by solving a least-squares problem:
@@
Finally, we select the stitching layer $k^{\star}$ that leads to the smallest (mean squared) error, $k^{\star}=\arg\min_{k}\|\mathbf{B}\mathbf{S}^{*}_{k}-\mathbf{A}_{k}\|_{F}^{2},$ and assemble the 3D VAE by concatenating $\mathcal{E}$, $\mathbf{S}^{*}_{k^{\star}}$ and $F_{k^{\star}+1:l}$. Empirically, we find that most combinations of foundational VAEs and 3D feedforward models can be stitched in this manner, with minimal performance loss.
@@ 3.2 Alignment via Direct Reward Finetuning
Alignment algorithm. To optimize the generative model according to the reward function above, we employ direct reward finetuning (Clark et al., 2023; Xu et al., 2023; Prabhudesai et al., 2024; Wu et al., 2024c; Shen et al., 2025). I.e., the model generates samples by unfolding the full denoising path, and the rewards computed from these samples are then backpropagated through the denoising chain. While the algorithm benefits from gradient-based feedback, it can also suffer from exploding gradient norms. To stabilize the optimization, we generalize the idea of DRTune (Wu et al., 2024c): gradients are detached from the inputs to the generative model, but retained during the update step to the next denoising state. In this way, reward propagation remains stable even at early denoising steps. Furthermore, we modify the optimizer for better computational efficiency by (i) randomized sampling, using fewer timesteps than during inference, and (ii) randomizing the subset of denoising steps where gradients are backpropagated, such that the model learns from diverse denoising trajectories. For further details, see Appendix B.2.

### v2

# Text-to-3D by Stitching a Multi-view Reconstruction Network to a Video Generator

@@ Abstract
The rapid progress of large, pretrained models for both visual content generation and 3D reconstruction opens up new possibilities for text-to-3D generation. Intuitively, one could obtain a formidable 3D scene generator if one were able to combine the power of a modern latent text-to-video model as “generator” with the geometric abilities of a recent (feedforward) 3D reconstruction system as “decoder”. We introduce VIST3A, a general framework that does just that, addressing two main challenges. First, the two components must be joined in a way that preserves the rich knowledge encoded in their weights. We revisit model stitching, i.e., we identify the layer in the 3D decoder that best matches the latent representation produced by the text-to-video generator and stitch the two parts together. That operation requires only a small dataset and no labels. Second, the text-to-video generator must be aligned with the stitched 3D decoder, to ensure that the generated latents are decodable into consistent, perceptually convincing 3D scene geometry. To that end, we adapt direct reward finetuning, a popular technique for human preference alignment. We evaluate the proposed VIST3A approach with different video generators and 3D reconstruction models. All tested pairings markedly improve over prior text-to-3D models that output Gaussian splats. Moreover, by choosing a suitable 3D base model, VIST3A also enables high-quality text-to-pointmap generation. footnotetext: Project page: https://gohyojun15.github.io/VIST3A/
@@ 1 Introduction
With image and video generators now a commodity, text-to-3D models that produce 3D scenes from text prompts have become a new research frontier, with applications in AR/VR, gaming, robotics, and simulation. Early methods for 3D generation adopt Score Distillation Sampling (SDS) (Poole et al., 2023; Tang et al., 2024b; Wang et al., 2023b; Chen et al., 2024b) to optimize a 3D representation, e.g. a NeRF (Mildenhall et al., 2021; Müller et al., 2022) or 3D Gaussian Splats (3DGS, Kerbl et al., 2023) under a pretrained 2D diffusion prior (Rombach et al., 2022). A drawback these methods have in common is the need for slow per-scene optimization. Another line of work uses multi-stage pipelines that first synthesize images and then lift them to 3D with a separate model (Tang et al., 2024a; Xu et al., 2024b; Zhang et al., 2024b) or with per-scene optimization (Gao et al., 2024; Wu et al., 2024a; Yu et al., 2025b); employ progressive warping and refinement (Shriram et al., 2025; Yu et al., 2025a; 2024); or sequentially chain multiple generative modules (Yang et al., 2025b; Engstler et al., 2025). The multi-stage design not only increases model complexity and engineering effort, but also makes such models prone to error accumulation (Lin et al., 2025; Meng et al., 2025).
@@
A recent trend is to directly generate the 3D representation with end-to-end latent diffusion models  (LDMs, Schwarz et al., 2025; LAN et al., 2025; Li et al., 2025b; a). A prominent line of work starts from pretrained 2D image (Esser et al., 2024; Rombach et al., 2022) or video models (Team, 2024; Yang et al., 2025e) and finetunes them to output multi-view 2D latents, reusing the pretrained priors (Szymanowicz et al., 2025; Liang et al., 2025; Schwarz et al., 2025; Lin et al., 2025; Yang et al., 2025c; Go et al., 2025a; b). Subsequently, a VAE-style decoder is trained to decode those latents into the desired 3D representation, see Fig. 2. The LDM-like design unifies 2D generation and multi-view reconstruction within the latent space and enables efficient 3D scene generation with a compact, well-amortized decoder.
@@
Still, two key limitations remain. First, we argue that the Achilles heel of existing 2D-to-3D diffusion models is the decoder. By simply repurposing the 2D VAE to produce 3D outputs, the network must learn 3D reconstruction more or less from scratch, which requires extensive training and large datasets that are hard to obtain (Yang et al., 2025c; Szymanowicz et al., 2025; Go et al., 2025b). This practice becomes increasingly problematic as new, better 3D foundation models emerge  (Wang et al., 2026; 2025b; 2024; Zhang et al., 2025) and the ad-hoc trained decoders of text-to-3D models fall further behind the state of the art in 3D vision.
@@
Second, the prevalent training scheme tends to suffer from weak alignment between the generative model and the VAE decoder. Typically, the former is finetuned on multi-view datasets with a generative objective like a diffusion loss (Song et al., 2021; Sohl-Dickstein et al., 2015; Ho et al., 2020) or flow matching (Liu et al., 2023; Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023), which only indirectly promotes 3D-consistent latents. Moreover, the separate training may cause the latents, even if 3D-consistent, to be out of domain from the perspective of the decoder. To mitigate that misalignment, it has been proposed to add rendering losses that promote decodable latents (Lin et al., 2025). However, the resulting objective is based on single-step sampling and does not sufficiently take into account the denoising trajectory, leading to weak alignment at inference.
@@
We introduce VIST3A: VIdeo VAE STitching and 3D Alignment. The proposed method consists of two complementary components that address the above-mentioned limitations, see Fig. 2. First, we resort to the concept of model stitching (Pan et al., 2023; Lenc and Vedaldi, 2015; Bansal et al., 2021; Csiszárik et al., 2021; Yang et al., 2022) to leverage powerful, pretrained feedforward 3D models for decoding, rather than start from scratch. The idea is to attach the relevant part of a 3D reconstruction network as a “decoder” to the latent space of a video VAE. For this to work, there needs to be one or more layers in the 3D model whose activations are similar (up to a linear transformation) to those in the VAE’s latent space, despite their independent pretraining. Perhaps surprisingly, this turns out to be the case. For the 3D model, we identify the layer with the most linear relation to the LDM latents, slice the network before that layer, and retain the downstream portion as 3D decoder. After fitting a single, linear stitching layer (in closed form), the VAE latent space already matches the expected input of the 3D decoder well, such that subsequent fine-tuning will be minor and not degrade the respective generative and 3D reasoning capabilities of the two base models.
@@
Second, we further improve alignment between the generative model and the stitched decoder through direct reward finetuning (Clark et al., 2024; Xu et al., 2023; Prabhudesai et al., 2024; Wu et al., 2024c; Shen et al., 2025). In that technique, commonly used to align diffusion models with human preferences, reward signals are defined based on the “goodness” of the VAE output – in our setting, the visual quality and 3D consistency of the decoded 3D representations. Maximizing these rewards encourages the LDM to produce latents that are 3D-consistent and lie within the decoder’s input domain, ensuring high-quality outputs. Importantly, our alignment compares video model outputs and images rendered from the generated 3D scenes, hence it does not require labels.
@@ 2 Related Works
3D generation. Recent works have explored various 3D representations for generative modelling, including point clouds (Mo et al., 2023; Nichol et al., 2022; Vahdat et al., 2022), meshes (Xu et al., 2024a), voxel grids (Sanghi et al., 2023), NeRFs (Chen et al., 2023; Müller et al., 2022; Mildenhall et al., 2021), and 3DGS (Henderson et al., 2024; Zhang et al., 2024a; Kerbl et al., 2023). Score distillation using 2D diffusion models is time-consuming, as it requires per-scene test time optimization (Wang et al., 2023a; Shi et al., 2024; Wang et al., 2023b), while multi-stage pipelines (Yu et al., 2025b; Liu et al., 2024; Zheng et al., 2025) lack robustness and create significant engineering overhead. For further details on multi-stage pipelines, please refer to Appendix A.
@@
To tackle such situations, recent methods (Szymanowicz et al., 2025; Liang et al., 2025; Schwarz et al., 2025; Lin et al., 2025; Yang et al., 2025c; Go et al., 2025a; b) repurpose the comprehensive knowledge of the visual world that is implicit in 2D image generators. The general strategy is to finetune a pretrained 2D model on multi-view data, by using generative losses to enforce cross-view consistency. In many cases training is further supported by additional 3D cues like camera poses (Li et al., 2024b; Go et al., 2025b), depthmaps (Go et al., 2025a; Yang et al., 2025c), or pointmaps (Szymanowicz et al., 2025). The resulting multi-view latents are decoded to 3D scenes with a dedicated VAE-style decoder, meaning that 3D reasoning capabilities must be rebuilt from scratch, and that they are only weakly aligned with the generator output – limitations which we address with VIST3A.
@@
Learned 3D reconstruction. A notable trend in 3D computer vision is the trend to move away from multi-stage pipelines and iterative optimization towards end-to-end, feedforward 3D modelling. Classical reconstruction pipelines based on SfM (Hartley and Zisserman, 2003; Schönberger and Frahm, 2016) and MVS (Furukawa et al., 2015; Schönberger et al., 2016) require incremental, iterative optimization, whereas recent advances like DUSt3R (Wang et al., 2024) and MASt3R (Leroy et al., 2024) directly predict 3D point maps in one forward pass. Several follow-up works have further reduced test-time optimization (Tang et al., 2025; Wang et al., 2025c; Yang et al., 2025a). Likewise, 3D Gaussian splatting has evolved from per-scene optimization to feedforward prediction (Charatan et al., 2024; Chen et al., 2024a; Ye et al., 2025). Once more, data scaling has been a critical factor (Wang et al., 2025b; 2026). Consequently, replicating the 3D capabilities of recent feedforward models as part of VAE training would be difficult and costly. VIST3A offers a solution by reusing, rather than rebuilding, models like AnySplat (Jiang et al., 2025), VGGT (Wang et al., 2025b), or MVDUSt3R (Tang et al., 2025).
@@
Model stitching. Recomposing the heads and tails of two different networks was initially studied as a way to assess the equivariance of neural representations (Lenc and Vedaldi, 2015), and as an experimental tool to compare two different representations (Csiszárik et al., 2021; Bansal et al., 2021). To ensure invariance against trivial affine transformations, the head of some trained network $A$ is normally attached to the tail of another network $B$ via a linear, trainable stitching layer. Besides revealing similarities between networks that common metrics like CKA (Kornblith et al., 2019) would miss, it was also found that different architectures that were trained on the same data can often be stitched into a new, hybrid model with minimal degradation (Bansal et al., 2021). This has opened the door for practical uses of stitching, e.g. DeRy (Yang et al., 2022) for resource-constrained reassembly of pretrained models and SN-Net (Pan et al., 2023) to build networks with varying scales. Going one step further, we demonstrate that strong 3D VAEs [^1] To be consistent with existing literature (LAN et al., 2025; Yang et al., 2025c), we also use the term “3D VAE”, although the mapping from 2D images to 3D scene is, technically, not a variational auto-encoder. can be obtained by stitching a foundational 3D model to the latent space of a video VAE as its decoder, even if they were trained independently on different data.
@@ 3.1 Model Stitching for 3D VAE Construction
Step 1: Finding the stitching index and initialization. To identify the layer $k^{*}$ in the 3D model whose representation is most compatible with the VAE latent, we first push a set of $N$ samples through the encoder $\mathcal{E}$ to obtain their latents $\mathbf{B}\in\mathbb{R}^{N\times D_{\mathcal{E}}}$. Here, $D_{\mathcal{E}}$ denotes the dimensionality of the encoder latent space, and $D_{F}^{k}$ denotes the dimensionality of the feature (activation) at layer $k$. Then, we scan over candidate layers $k\in\{1,...,l-1\}$ of the 3D model and, for each layer in turn, extract the activations $\mathbf{A}_{k}\in\mathbb{R}^{N\times D_{F}^{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}k}}}$ and fit the linear stitching layer $\mathbf{S^{*}}_{k}\in\mathbb{R}^{D_{\mathcal{E}}\times D_{F}^{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}k}}}$ that best recovers the activations of the 3D model at layer $k$, by solving a least-squares problem:
@@
Finally, we select the stitching layer $k^{\star}$ that leads to the smallest (mean squared) error, $k^{\star}=\arg\min_{k}\|\mathbf{B}\mathbf{S}^{*}_{k}-\mathbf{A}_{k}\|_{F}^{2},$ and assemble the 3D VAE by concatenating $\mathcal{E}$, $\mathbf{S}^{*}_{k^{\star}}$ and $F_{k^{\star}+1:l}$. Empirically, we find that many combinations of foundational VAEs and 3D feedforward models can be stitched in this manner, with minimal performance loss.
@@ 3.2 Alignment via Direct Reward Finetuning
Alignment algorithm. To optimize the generative model according to the reward function above, we employ direct reward finetuning (Clark et al., 2024; Xu et al., 2023; Prabhudesai et al., 2024; Wu et al., 2024c; Shen et al., 2025). I.e., the model generates samples by unfolding the full denoising path, and the rewards computed from these samples are then backpropagated through the denoising chain. While the algorithm benefits from gradient-based feedback, it can also suffer from exploding gradient norms. To stabilize the optimization, we generalize the idea of DRTune (Wu et al., 2024c): gradients are detached from the inputs to the generative model, but retained during the update step to the next denoising state. In this way, reward propagation remains stable even at early denoising steps. Furthermore, we modify the optimizer for better computational efficiency by (i) randomized sampling, using fewer timesteps than during inference, and (ii) randomizing the subset of denoising steps where gradients are backpropagated, such that the model learns from diverse denoising trajectories. For further details, see Appendix B.2.